# Deep Sub-Ensembles Meets Quantile Regression: Uncertainty-aware Imputation for Time Series

### v3

@@ Abstract
Real-world time series data frequently have significant amounts of missing values, posing challenges for advanced analysis. A common approach to address this issue is imputation, where the primary challenge lies in determining the appropriate values to fill in. While previous deep learning methods have proven effective for time series imputation, they often produce overconfident imputations, which could brings a potentially overlooked risk to the reliability of the intelligent system. Diffusion methods are proficient in estimating probability distributions but face challenges with high missing rates and moreover, computationally expensive due to the nature of the generative model framework. In this paper, we propose Quantile Sub-Ensembles, a novel method to estimate uncertainty with ensemble of quantile-regression-based task networks and then incorporate Quantile Sub-Ensembles into a non-generative time series imputation method. Our method not only produces accurate imputations that is robust to high missing rates, but also is computationally efficient due to the fast training of its non-generative model. We examine the performance of the proposed method on two real-world datasets, the air quality and health-care datasets, and conduct extensive experiments to show that our method outperforms other most of the baseline methods in making deterministic and probabilistic imputations. Compared with the diffusion method, CSDI, our approach can obtain comparable forecasting results which is better when more data is missing, and moreover consumes a much smaller computation overhead, yielding much faster training and test.
@@ 1 Introduction
Multivariate time series are used in a variety of real-world applications, including meteorology [37; 40], financial marketing [3; 4], health care [11; 32], load forecasting [21] and artificial intelligence for IT Operations (AIOps) [31]. Time series data serve as prevalent signals for classification and regression tasks across various applications. However, missing values commonly appear in time series due to data corruptions, merging irregularly sampled data and human recording errors. This has been recognized as a severe problem in time series analysis and downstream applications [41; 50; 44].
@@
The traditional time series imputation methods are divided into two categories. One is deletion, which removes partially observed samples and features, resulting in biased parameter estimates [25]. The other is the statistical method, where missing values are imputed using median or average values, but this method is unsuitable for high-precision scenarios. Consequently, numerous studies have resorted to machine-learning approaches to address the assignment of imputing missing values. However, some studies show that traditional machine-learning methods such as K-Nearest Neighbor (KNN) and Multivariate Imputation by Chained Equations (MICE) [2] are not optimal for large-scale data.
@@
In recent years, deep neural networks have achieved remarkable success for time series imputation. Recurrent Neural Networks (RNNs) based models can achieve high accuracy by capturing temporal dependencies of time series in imputing missing values [42; 10]. Nevertheless, they can not quantify uncertainty and derive a reliable imputation rule [27]. To address this issue, Bayesian Neural Networks (BNNs) estimate uncertainty in their parameter space by integrating Bayesian probability theory to impute missing values [43]. However, performing Bayesian inference over the high-dimensional network weights is computationally expensive and challenging. Recently, the score-based diffusion model such as CSDI [45] has been proposed to learn the conditional distribution for probabilistic imputation. Diffusion models primarily rely on the generative distribution learned from complete data, which may lead to degenerative results when handing data with high missing rate. This occurs because the model may lack sufficient known data points to accurately infer the distribution characteristics of the missing data, leading to increased imputation errors [9]. In addition, the probabilistic inference of the generative diffusion model has a high computational complexity which results in slow training convergence and large computing overhead.
@@
In view of this, we introduce Quantile Sub-Ensembles, which produces high quality uncertainty estimates with ensembling quantile-regression-based task networks of the model, while sharing a common trunk network. It offers a robust solution to the phenomenon of overfitting to the mean and extreme values. Subsequently, we equip bidirectional long short-term memory networks (BiLSTM) with Quantile Sub-Ensembles to quantify inherent uncertainty while making accurate imputation. Empirically, we conduct extensive experiments on two real-world datasets: air quality and health-care to evaluate the performance of the proposed method, which shows that our method outperforms all other models at 90% missing rate and is only slightly inferior to CSDI at 10% and 50% missing rate in terms of imputation performance and uncertainty quantification. Moreover, the proposed method consumes a much smaller computation overhead, yielding much faster training and test speed than CSDI.
@@
    We incorporate Quantile Sub-Ensembles into BiLSTM for probabilistic time series imputation, which can quantify uncertainty during the imputation process.
@@
    We evaluate the proposed method on two representative datasets: health-care and air quality, which shows our method not only achieves high imputation accuracy but also produces reliable uncertainty especially in a higher missing rate. Moreover, it is significantly more computationally efficient than diffusion models, and consumes much less training and test time.
@@ 2.1 Statistical methods
Over the past few decades, numerous imputation methods have been employed to address missing values in multivariate time series. Statistical methods are widely used for time series imputation due to their user-friendly nature and computational efficiency. For example, since there is a strong tendency for time series, we can impute missing values through the trend of the fitting data changes, which is called linear interpolation [5]. Because of their simplicity, these imputation methods are often inadequate in situations that demand high precision. The autoregressive methods such as ARIMA [1] are applied to fill the missing values. It converts the data into stationary data through differential analysis, and then regresses the dependent variable only to its lagged value, as well as the present value and lagged value of the random error term. These methods often struggle to effectively impute missing values due to their neglect of dependencies between time points and the relationship between observed and imputed values.
2.2 Traditional machine-learning based approaches
Traditional machine-learning based approaches are widely applied to time series imputation. For instance, linear regression models impute missing values by leveraging the linear relationship between the target variable and one or more input features. Lagged values or relevant features can be used as inputs for imputation. The K-Nearest Neighbors (KNN) algorithm [23] can also be applied in the time series imputation to estimate missing values by finding the k-nearest data points in terms of similarity within the time series. While KNN is a straightforward and interpretable imputation method for time series data, it has limitations related to parameter sensitivity, computational complexity and its ability to handle various data characteristics effectively. Multivariate Imputation by Chained Equations (MICE) [2] initializes the missing values arbitrarily and then each missing variable is estimated according to the chain equations, which may not be suitable for large-scale data.
@@ 2.3 Deep learning models
Recently, deep learning models have demonstrated powerful learning capabilities on multivariate time series imputation. Recurrent Neural Networks(RNNs) models [51; 10; 39] have been optimized to capture the temporal dependencies and do not impose specific assumptions. For example, BRITS [10] imputes missing values according to hidden states from BRNNs, which learns the temporal dependencies from time series directly. The self-attention-based model, SAITS [18], learns missing values by a joint-optimization training approach of imputation and reconstruction. Moreover, the self-attention mechanism is incorporated into RNNs [42] to capture global and local information from time series and distant dependencies. While these methods have high accuracy, this is not sufficient to characterize the performance in time series imputation where uncertainty factors can influence the performance tremendously.
Bayesian Neural Networks (BNNs) are particularly well-suited for uncertainty estimation [43]. Unlike traditional neural networks, BNNs offer probabilistic estimations for their weights and biases, making them valuable for tasks that require uncertainty quantification and robust predictions. However, an incorrect prior choice may lead to poor model performance or difficulty in convergence [6]. Denoising diffusion probabilistic models [22; 45] have surpassed existing models in many deep learning tasks. As an innovative imputation approach, they offer probabilistic estimation and learn the conditional distribution using conditional score-based diffusion models. However, there are instances when its benefits can also translate into substantial drawbacks. The diffusion models for the imputation task may face challenges when dealing data with high missing rate. Because the generative model lacks sufficient data points to infer the distribution characteristics of the missing data [9]. In addition, the high computational complexity of the probabilistic inference process in conditional score-based diffusion models leads to slow training convergence and substantial consumption of computing resources.
@@ 2.4 Quantile regression
Quantile regression techniques [17; 46; 36] have been introduced to learn the probability distribution of state-action values, facilitating the modeling of risks across various confidence intervals and fortifying resilience against extreme scenarios, respectively. This approach, shared across the referenced works, revolves around the pivotal step of apportioning layers among different quantiles, a methodology also embraced within the framework of this study. Quantile regression is less sensitive to outliers compared to ordinary least squares (OLS) regression [15]. It provides a more reliable estimation with using the conditional quantiles compared to OLS, making it suitable for data with extreme values or skewness. In traditional quantile regression, the assumption is that of constant variance. However, in our model, we observe that the variance under different quantiles varies with the input variables. This leads us to adopt a heteroscedasticity assumption, which allows for a more flexible and accurate modeling of the data [19]. Based on these advantages, quantile regression is widely used in economics and finance to analyze income distribution [30], asset pricing [24] and risk assessment [13]. By employing quantile regression, our model accounts for the initial distribution of the data when filling in missing values. This leads to imputations that vary within a narrower range, thereby reducing the dispersion of the imputed results. This method effectively tackles the inherent uncertainty present in the data. In this scenario, each quantile represents a specific percentile, such as the median (50th percentile) or the 90th percentile. The width of the estimated quantile (i.e., prediction intervals) can provide valuable information about the uncertainty associated with the predictions [16].
3 Preliminaries
@@
Figure 1: A conceptual comparison between Deep Ensembles and Deep Sub-Ensembles with $n$ ensemble members is illustrated. In (a), multiple trunk networks, denoted as $\mathcal{F}_{i}$ are independently trained on sample data $\mathcal{S}_{i}$ from dataset $\mathcal{D}$. In contrast, (b) shows that a single trunk network, $\mathcal{F}_{share}$, is shared across all ensemble members. $\mathcal{K}_{i}$ represent task networks with $n$ ensemble members. Both methods involve combining ensemble predictions to generate outputs that reflect uncertainty.
@@
Deep Ensembles and Deep Sub-Ensembles. Deep Ensembles [29] is a straightforward yet highly effective approach for quantifying predictive uncertainty. Deep Ensembles consists of multiple trunk networks that are trained independently on sample data and combines ensemble predictions by uniform voting which is shown in Figure 1(a). Deep Ensembles is widely used to improve model robustness and performance [49] by alleviating the over-fitting of DNNs and capturing different patterns in the data, ultimately enhancing the reliability and accuracy of predictions. Performing Bayesian inference over DNNs weights (i.e., learning weight distributions) is very challenging in highly nonlinear and high-dimensional space of DNNs. Deep Ensembles directly combines the outputs of different DNNs and are a prevalent tool for producing uncertainty, which is well-suited for distributed computing environments and makes it particularly appealing for large-scale deep learning applications owing to the simple implementation. Besides, Deep Ensembles offer a valuable means of estimating predictive uncertainty stemming from neural networks and effectively calibrating unidentified classes within supervised learning problems [29].
@@
Due to the parallel training and inference across multiple neural network models with the same structure in Deep Ensembles, the computational overhead becomes enormous as ensemble members increase. Deep Sub-Ensembles [47] is motivated by the observation that deep neural networks inherently learn a hierarchical structure of features, where the complexity of features increases with depth, culminating in the final layers that execute specific tasks like classification or regression. In this case, a neural network can be separated into two sub-networks, the trunk network $\mathcal{F}$ which tends to learn similar features across different ensemble members, and the task network $\mathcal{K}$ which play a more significant role in contributing to uncertainty. The objective of Deep Sub-Ensembles is to create an ensemble comprising a shared trunk network $\mathcal{F}_{share}$ and multiple instances of task-specific networks $\mathcal{K}_{i}$ to form a sub-ensemble. The architecture is shown in Figure 1(b).
4 Methodology
In this section, we first present the problem formulation and introduce how to incorporate quantile regression into Deep Sub-Ensembles, which we refer to as Quantile Sub-Ensembles. We then apply Quantile Sub-Ensembles to capture uncertainty in time series imputation, proposing to enhance the BiLSTM model, a representative time series imputation approach with Quantile Sub-Ensembles to efficiently quantify uncertainty. Finally, we provide theoretical discussions for this method.
4.1 Multivariate time series imputation
We denote a multivariate time series as $X=\{x_{1:T}^{1:K}\}\in\mathbb{R}^{T\times K}$ where $T$ represents the length of time series and $K$ is the dimension of features. For example, $x_{t}^{k}$ represents the $k$-th feature value of $x$ at the timestamp $d_{t}$. In the real-world setting, a time series may have some missing values due to unexpected accidents. Therefore, in order to train conveniently, We define a mask vector $M=\{m_{1:T}^{1:K}\}\in\mathbb{R}^{T\times K}$ to represent the missing values in $X$ where $m_{t}^{k}=0$ if $x_{t}^{k}$ is not observed, otherwise $m_{t}^{k}=1$.
$\delta_{t}^{k}=\begin{cases}0,&\textrm{if}\quad t=1,\\ d_{t}-d_{t-1}+\delta_{t-1}^{k},&\textrm{if}\quad t>1\quad\textrm{and}\quad m_{t-1}^{k}=0,\\ d_{t}-d_{t-1},&\textrm{if}\quad t>1\quad\textrm{and}\quad m_{t-1}^{k}=1.\end{cases}$
In the real-world setting, we need to impute the missing values in $X$ and recover the entire time series, which is called time series imputation. In this paper, we focus on probabilistic time series imputation [20], i.e., estimating the probabilistic distribution of time series missing values through observation. Therefore, uncertainty can be quantified while imputing missing values.
@@ 4.2 Quantile Sub-Ensembles
In time series, values can be missing for either continuous segments or sporadic time points. Such random missing patterns introduce significant uncertainty into our imputation results, which can adversely affect the performance of downstream tasks. We develop Quantile Sub-Ensembles, an algorithm to quantify uncertainty with training multiple quantile-regression-based task networks and alleviate overfitting to the mean sample intensity and extreme values. The concept is shown in Figure 2.
Figure 2: The algorithm of Quantile Sub-Ensembles. We set $N$ task networks $\{\mathcal{K}_{1},\mathcal{K}_{2},...,\mathcal{K}_{N}\}$ with the same architecture corresponding to $N$ ensemble members. In particular, task networks are defined by the quantiles $\{q_{1},q_{2},...,q_{N}]\}$. We stack the trunk model $\mathcal{F}_{share}$ and each task network $\mathcal{K}_{i}$ with randomly initialized parameters $\theta_{i}$ and train an instance of it. The output from $\mathcal{K}_{i}$ will be passed into quantile regression loss function $L$ to calculate the loss $l_{i}$ with the corresponding quantile $q_{i}$.
Quantile regression [28] is a statistical technique that extends traditional linear regression by modeling different quantiles of the response variable’s distribution. While linear regression focuses solely on estimating the conditional mean, quantile regression aims to estimate the conditional quantiles (e.g., the 25th, 50th, or 75th percentiles). This approach provides a more nuanced understanding of the relationship between predictors and the response variable, offering insights across various points of the distribution rather than just the mean.
Quantile regression aims to estimate a specific quantile of the response variable. The conditional distribution function [38] of $Y$ given $X=x$ is:
@@
$l(x,y,q_{i})=\begin{cases}q_{i}(x-y)&if\quad x\geq y\\ (1-q_{i})(y-x)&if\quad x<y\end{cases}$ (3)
@@
We consider the ensemble as a mixture model [29] with uniform weighting across its components and combine the outputs as $p(y|x)=N^{-1}\sum_{1}^{N}p_{\theta_{i}}(y|x,\theta_{i})$. For regression tasks, the output is modeled as a mixture of Gaussian model with all distributions equally weighted. To simplify the computation of quantiles and predictive probabilities, we approximate the ensemble output with a Gaussian distribution whose mean $u_{\theta_{i}}(x)$ and variance $\sigma_{\theta_{i}}^{2}(x)$ correspond to the mean $u_{*}(x)$ and variance $\sigma_{*}^{2}(x)$ of the mixture:
@@
Regression necessitates a distinct loss function, as supervision is available for $u_{\theta_{i}}(x)$, but not for $\sigma_{\theta_{i}}^{2}(x)$. To address this, we integrate a heteroscedastic Gaussian log-likelihood loss [26] [29] into quantile regression loss:
$-logp(y|x)=\frac{log\sigma_{\theta_{i}}^{2}(x)}{2}+\frac{l_{i}(u_{\theta_{i}}(x),y,q_{i})}{2\sigma_{\theta_{i}}^{2}(x)}.$ (7)
@@
With this loss function, the output variance $\sigma_{\theta_{i}}^{2}(x)$ can be interpreted as an estimate of the aleatoric uncertainty in the data, while epistemic uncertainty is captured through ensembling and is represented by $\sigma_{*}^{2}(x)$ [47].

### v4

@@ Abstract
Real-world time series data often exhibits substantial missing values, posing challenges for advanced analysis. A common approach to addressing this issue is imputation, where the primary challenge lies in determining the appropriate values to fill in. While previous deep learning methods have proven effective for time series imputation, they often produce overconfident imputations, which poses a potentially overlooked risk to the reliability of the intelligent system. Diffusion methods are proficient in estimating probability distributions but face challenges under a high missing rate and are, moreover, computationally expensive due to the nature of the generative model framework. In this paper, we propose Quantile Sub-Ensembles, a novel method that estimates uncertainty with ensembles of quantile-regression-based task networks and incorporate Quantile Sub-Ensembles into a non-generative time series imputation method. Our method not only produces accurate and reliable imputations, but also remains computationally efficient due to its non-generative framework. We conduct extensive experiments on five real-world datasets, and the results demonstrates superior performance in both deterministic and probabilistic imputation compared to baselines across most experimental settings. The code is available at https://github.com/yingliu-coder/QSE.
keywords:
Time series imputation, Uncertainty quantification, Deep ensembles, Quantile regression
@@ 1 Introduction
Multivariate time series are used in a variety of real-world applications, including meteorology [weather2], financial marketing [financial2], material science [gao2024dilated], and artificial intelligence for IT Operations (AIOps) [li2022stackvae]. Time series data serves as prevalent signals for classification and regression tasks across various applications. However, missing values commonly appear in time series due to data corruptions, merging irregularly sampled data, and human operational errors. This has been recognized as a severe problem in time series analysis and downstream applications.
@@
The traditional time series imputation methods are divided into two categories. One is deletion, which removes partially observed samples and features, resulting in biased parameter estimates. The other is the statistical method, where missing values are imputed using median or average values. However, this approach is unsuitable for high-precision scenarios. Consequently, numerous studies have resorted to machine-learning approaches to address the assignment of imputing missing values. However, some studies show that traditional machine-learning methods such as K-Nearest Neighbor (KNN) [KNN], Multivariate Imputation by Chained Equations (MICE) [MICE], and Matrix Decomposition [matricdec] are not optimal for large-scale data.
@@
In recent years, deep neural networks have achieved remarkable success in time series imputation. RNN-based models can achieve high accuracy by capturing temporal dependencies of time series in imputing missing values [BRITS, TKAE]. Nevertheless, they cannot quantify uncertainty and derive a reliable imputation rule. To address this issue, Bayesian Neural Networks (BNNs) estimate uncertainty in their parameter space by integrating Bayesian probability theory to impute missing values [BNNI]. However, performing Bayesian inference over the high-dimensional network weights is computationally expensive and challenging. Recently, the score-based diffusion model such as CSDI [CSDI] has been proposed to learn the conditional distribution for probabilistic imputation. Diffusion models primarily rely on the generative distribution learned from the observations, which may lead to degenerative results when handling data with a high missing rate [cao2024survey]. In addition, the generative diffusion model entails sampling-intensive Markov chain iterations which fundamentally contribute to the prohibitive computational costs during the inference process.
@@
In view of this, we introduce Quantile Sub-Ensembles, which produces high-quality uncertainty estimates with ensembling quantile-regression-based task networks of the model, while sharing a common trunk network. It offers a robust solution to the phenomenon of overfitting to the mean and extreme values. Subsequently, we equip bidirectional long short-term memory networks (BiLSTM) with Quantile Sub-Ensembles to quantify inherent uncertainty while making accurate imputation. Empirically, we conduct extensive experiments on five real-world datasets (solar, traffic, healthcare, electricity, and air-quality) to evaluate the performance of the proposed method, which shows that our method outperforms the baseline models on the most evaluation metrics in terms of imputation accuracy and uncertainty quantification. Moreover, the proposed method delivers substantial computational advantages with significantly faster training and inference speeds compared to generative diffusion models.
@@
    We incorporate Quantile Sub-Ensembles into BiLSTM for probabilistic time series imputation, which can estimate the predictive uncertainty effectively during the imputation process.
@@
    We evaluate the proposed method on five representative datasets, which shows our method not only achieves high imputation accuracy but also produces reliable uncertainty especially in a higher missing rate. Moreover, it is significantly more computationally efficient than diffusion models.
@@ 2.1 Statistical methods
Over the past few decades, numerous imputation methods have been employed to address missing values in multivariate time series. Statistical methods are widely used for time series imputation due to their user-friendly nature and computational efficiency. For example, since there is a strong tendency for time series, we can impute missing values through the trend of the fitted data changes, which is called linear interpolation [linearInterpolation]. Because of their simplicity, these imputation methods are often inadequate in situations that demand high precision. Autoregressive methods such as ARIMA [ARIMA] are applied to fill the missing values. It converts the data into stationary data through differential analysis, and then regresses the dependent variable only on its lagged value, as well as the present value and lagged value of the random error term. These methods often struggle to effectively impute missing values due to their neglect in capturing dependencies across time points and the relationship between observed and imputed values.
2.2 Traditional machine-learning-based approaches
Traditional machine-learning-based approaches are widely applied to time series imputation. For instance, linear regression models impute missing values by leveraging the linear relationship between the target variable and one or more input features. Lagged values or relevant features can be used as inputs for imputation. The K-Nearest Neighbors (KNN) algorithm [KNN] can also be applied to the time series imputation such as MedImpute [bertsimas2021imputation], which imputes missing data by integrating patient-specific temporal information into a flexible framework. While KNN is a straightforward and interpretable imputation method for time series data, it has limitations related to parameter sensitivity, computational complexity and its ability to handle various data characteristics effectively. Multivariate Imputation by Chained Equations (MICE) [MICE] initializes the missing values arbitrarily and each missing variable is estimated according to the chain equations, which may not be suitable for large-scale data.
@@ 2.3 Deep learning models
Recently, deep learning models have demonstrated powerful learning capabilities on multivariate time series imputation. Recurrent Neural Networks (RNNs) have been optimized to capture the temporal dependencies and do not impose specific assumptions. For example, BRITS [BRITS] imputes missing values according to hidden states from BRNNs, which learns the temporal dependencies from time series directly. TKAE [TKAE] generates compressed representations of multivariate time series through a recurrent neural network-based autoencoder to effectively handle missing data. The self-attention-based model, SAITS [SAITS], learns missing values by a joint-optimization training approach of imputation and reconstruction. Moreover, the self-attention mechanism is incorporated into RNNs [GLIMA] to capture global and local information from time series and distant dependencies. While these methods have high accuracy, this is not sufficient to characterize the performance in time series imputation where uncertainty factors can influence the performance tremendously.
Bayesian Neural Networks (BNNs) are particularly well-suited for uncertainty estimation [BNNI]. Unlike traditional neural networks, BNNs provide probabilistic distributions over their weights and biases, making them valuable for tasks that require uncertainty quantification and robust predictions. However, an incorrect prior choice may lead to poor model performance or difficulty in convergence [bayesianprior]. Denoising diffusion probabilistic models [Diffusion, CSDI, Score-CDM] have surpassed existing models in many deep learning tasks. As an innovative imputation approach, they offer probabilistic estimation and learn the conditional distribution using conditional score-based diffusion models. However, there are instances when their benefits can also translate into their drawbacks. Diffusion models for imputation may face challenges when handling data with a high missing rate, as the generative model lacks sufficient data points to infer the distributional characteristics of missing values [cao2024survey]. In addition, the high computational complexity of the probabilistic inference process in conditional score-based diffusion models leads to slow training convergence and substantial consumption of computing resources.
@@ 2.4 Quantile regression
Quantile regression techniques [QR-DQN] have been introduced to learn the probability distribution of state-action values, facilitating the modeling of risks across various confidence intervals and fortifying resilience against extreme scenarios. This approach, shared across the referenced works, revolves around the pivotal step of apportioning layers among different quantiles, a methodology also embraced within the framework of this study. Quantile regression is less sensitive to outliers compared to ordinary least squares (OLS) regression [yadav2024exponential]. It provides a more reliable estimation by using the conditional quantiles compared to OLS, making it suitable for data with extreme values or skewness characteristics. In traditional quantile regression, the model assumes constant variance. However, in our model, we observe that the variance under different quantiles varies with the input variables. This leads us to adopt a heteroscedasticity assumption, which allows for a more flexible and accurate modeling of the data [Heteroscedasticity]. Based on these advantages, quantile regression is widely used in economics and finance to analyze income distribution [income], asset pricing [pricing] and risk assessment [risk]. By employing quantile regression, our model accounts for the initial distribution of the data when filling in missing values. This narrows the range of imputed values, reducing their dispersion. This method effectively tackles the inherent uncertainty present in the data. In this scenario, each quantile represents a specific percentile, such as the median (50th percentile) or the 90th percentile. The width of the estimated quantile (i.e., prediction intervals) can provide valuable information about the uncertainty associated with the predictions [calibration].
3 Methodology
In this section, we first establish the theoretical foundation by analyzing uncertainty quantification mechanisms in Deep Ensembles and their parameter-efficient variant Deep Sub-Ensembles. Subsequently, we present the integration of quantile regression with Deep Sub-Ensembles, yielding the proposed Quantile Sub-Ensembles that generates uncertainty estimates through parallel task networks. We then implement this approach by enhancing the BiLSTM architecture with Quantile Sub-Ensembles for uncertainty-aware time series imputation. Finally, we provide theoretical discussions for this method.
@@ 3 Methodology
Figure 1: A conceptual comparison between Deep Ensembles and Deep Sub-Ensembles with NN ensemble members is illustrated. In (a), multiple trunk networks, denoted as $\mathcal{F}_{i}$ are independently trained on sample data $\mathcal{S}_{i}$ from dataset $\mathcal{D}$. In contrast, (b) shows that a single trunk network, $\mathcal{F}_{share}$, is shared across all ensemble members. $\mathcal{K}_{i}$ represent task networks with $n$ ensemble members. Both methods involve combining ensemble predictions to generate outputs that reflect uncertainty.
3.1 Deep Ensembles and Deep Sub-Ensembles
@@
Deep Ensembles [deepEnsembles] is a straightforward yet highly effective approach for quantifying predictive uncertainty. Deep Ensembles consists of multiple trunk networks that are trained independently on sample data and combines ensemble predictions by uniform voting which is shown in Fig. 1(a). Deep Ensembles is widely used to improve model robustness and performance [de1] by alleviating the over-fitting of DNNs and capturing different patterns in the data, ultimately enhancing the reliability and accuracy of predictions. Performing Bayesian inference over DNNs weights (i.e., learning weight distributions) is very challenging in highly nonlinear and high-dimensional space of DNNs. Deep Ensembles directly combines the outputs of different DNNs and are a prevalent tool for producing uncertainty, which is well-suited for distributed computing environments and makes it particularly appealing for large-scale deep learning applications owing to the simple implementation. Besides, Deep Ensembles offer a valuable means of estimating predictive uncertainty stemming from neural networks and effectively calibrating unidentified classes within supervised learning problems [deepEnsembles].
@@
Due to the parallel training and inference across multiple neural network models with the same structure in Deep Ensembles, the computational overhead becomes enormous as ensemble members increase. Deep Sub-Ensembles [sub-ensembles] is motivated by the observation that deep neural networks inherently learn a hierarchical structure of features, where the complexity of features increases with depth, culminating in the final layers that execute specific tasks like classification or regression. In this case, a neural network can be separated into two sub-networks, the trunk network $\mathcal{F}_{share}$ which tends to learn similar features across different ensemble members, and the task network $\mathcal{K}$ which play a more significant role in contributing to uncertainty. The objective of Deep Sub-Ensembles is to create an ensemble comprising a shared trunk network $\mathcal{F}_{share}$ and multiple instances of task-specific networks $\mathcal{K}_{i}$ to form a sub-ensemble. The architecture is shown in Fig. 1(b). In this paper, we adopt Deep Sub-Ensembles, which only ensembles the task network while sharing the trunk network. This method still produces high quality uncertainty estimates, yet significantly accelerates inference, since the trunk network requires only one forward pass and the smaller task network requires multiple but lightweight passes [sub-ensembles]. Consequently, the model efficiency is improved.
@@ 3.2 Quantile Sub-Ensembles
In time series, values can be missing for either continuous segments or sporadic time points. Such random missing patterns introduce significant uncertainty into our imputation results, which can adversely affect the performance of downstream tasks. We develop Quantile Sub-Ensembles, an algorithm to quantify uncertainty with training multiple quantile-regression-based task networks and alleviate overfitting to the mean sample intensity and extreme values. The concept is shown in Fig. 2.
Figure 2: The framework of Quantile Sub-Ensembles. We stack the trunk network $\mathcal{F}_{share}$ and NN parallel task networks $\{\mathcal{K}_{1},\mathcal{K}_{2},...,\mathcal{K}_{N}\}$ with the same architecture corresponding to NN ensemble members. The trunk network $\mathcal{F}_{share}$ functions as a shared feature extractor that learns temporal patterns from the input time series. This design not only reduces redundant parameterization but also improves efficiency. Subsequently, the task networks learn quantiles $\{q_{1},q_{2},...,q_{N}\}$ through randomly initialized parameters $\{\theta_{1},\theta_{2},...,\theta_{N}\}$. The imputations of all quantiles will be combined to obtain the reliable result.
Quantile regression [QR] is a statistical technique that extends traditional linear regression by modeling different quantiles of the response variable’s distribution. While linear regression focuses solely on estimating the conditional mean, quantile regression aims to estimate the conditional quantiles (e.g., the 10th or 90th percentiles). This approach provides a more nuanced understanding of the relationship between predictors and the response variable, offering insights across various points of the distribution rather than just the mean.
Quantile regression aims to estimate a specific quantile of the response variable. The conditional distribution function [CQR] of $Y$ given $X=x$ is:
@@
$l(x,y,q_{i})=\begin{cases}q_{i}(x-y)&if\quad x\geq y\\ (1-q_{i})(y-x)&if\quad x<y\end{cases},$ (3)
@@
We consider the ensemble as a mixture model [deepEnsembles, mikalsen2018time] with uniform weighting across its components and combine the outputs as $p(y|x)=N^{-1}\sum_{i=1}^{N}p_{\theta_{i}}(y|x,\theta_{i})$. For regression tasks, the output is modeled as a mixture of Gaussian model with quantiles are derived from the cumulative distribution of a single fitted distribution. Therefore, our method naturally guarantees ordered quantiles without the risk of crossing. To simplify the computation of quantiles and predictive probabilities, we approximate the ensemble output with a Gaussian distribution whose mean $u_{\theta_{i}}(x)$ and variance $\sigma_{\theta_{i}}^{2}(x)$ correspond to the mean $u_{*}(x)$ and variance $\sigma_{*}^{2}(x)$ of the mixture:
@@
Regression necessitates a distinct loss function, as supervision is available for $u_{*}(x)$, but not for $\sigma_{*}^{2}(x)$. To address this, we integrate a heteroscedastic Gaussian log-likelihood loss [uncertaintiesneed, deepEnsembles] into quantile regression loss:
$logp(y|x)=\frac{log\sigma_{*}^{2}(x)}{2}+\frac{N^{-1}\sum_{i=1}^{N}{l_{i}(u_{\theta_{i}}(x),y,q_{i})}}{2\sigma_{*}^{2}(x)}.$ (7)
@@
With this loss function, the output variance $\sigma_{*}^{2}(x)$ represents the predictive uncertainty estimate.